# compat=1.1.0

### v1

@@ Abstract
Dark matter (DM), in spite of being stable or long-lived on cosmological scales, can decay in the early Universe due to finite-temperature effects. In particular, a first order phase transition (FOPT) in the early Universe can provide a finite window for such decay, guaranteeing DM stability at lower temperatures, consistent with observations. The FOPT can lead to the generation of stochastic gravitational waves (GW) with peak frequencies correlated with DM mass. On the other hand, early DM decay into neutrinos can create a large neutrino asymmetry which can have interesting cosmological consequences in terms of enhanced effective relativistic degrees of freedom $N_{\rm eff}$, providing a solution to the recently observed Helium anomaly among others. Allowing DM decay to occur below sphaleron decoupling temperature, thereby avoiding overproduction of baryon asymmetry, forces the FOPT to occur at sub-electroweak scale. This leaves the stochastic GW within range of experiments like LISA, $\mu$ARES, NANOGrav etc.
@@ I Introduction
If a mechanism to explain BAU also leaves a large neutrino asymmetry at low temperatures, it can have interesting implications for observations related to BBN and CMB. Although the observed BAU restricts the net lepton asymmetry around sphaleron decoupling temperature $(T_{\rm sph})$ to be of same order as $\eta_{B}$, it is possible to have a large lepton asymmetry at post-sphaleron epoch $(T<T_{\rm sph})$ if the asymmetry is stored in neutrinos [^1] Charge neutrality of the early Universe restricts the charge lepton asymmetry to be $\leq\eta_{B}$.. Such large neutrino asymmetry can enhance the effective relativistic degrees of freedom $N_{\rm eff}$ which can be measured at CMB experiments Escudero:2022okz . Additionally, a large neutrino asymmetry can also alter the predictions of BBN with observable consequences Mangano:2010ei ; Mangano:2011ip ; Castorina:2012md ; Escudero:2022okz . A large neutrino asymmetry also allows resonant production of sterile neutrino DM via Shi-Fuller mechanism Shi:1998km  [^2] One may refer to a review Drewes:2016upu for details of these mechanisms and relevant experimental constraints.. Alternately, such asymmetries can also alter the cosmological bounds on light sterile neutrinos Mirizzi:2012we . Large neutrino asymmetry can also have other consequences for gravitational wave, baryogenesis, Barrie:2017mmr ; Gao:2024fhm , astrophysical structure formation Zeng:2018pcv etc.
@@ II The framework
$-\mathcal{L}\supset\big{(}y_{1}\overline{\chi}\Phi_{2}\nu+{\rm h.c.}\big{)}+V(\Phi_{1},\Phi_{2})$ (2)
@@
$V_{\rm tree}\equiv V(\Phi_{1},\Phi_{2})$ $=\lambda_{1}\left(|\Phi_{1}|^{2}-\frac{v_{D}^{2}}{2}\right)^{2}+\mu_{\Phi_{2}}^{2}|\Phi_{2}|^{2}+\lambda_{2}|\Phi_{2}|^{4}+\lambda_{\Phi_{1}\Phi_{2}}|\Phi_{1}|^{2}|\Phi_{2}|^{2}.$ (3)
In the above scalar potential $v_{D}$ is the vacuum expectation value (VEV) acquired by the scalar field $\Phi_{1}$ while driving the FOPT. A gauged $U(1)_{D}$ symmetry is considered under which $\Phi_{1}$ has charge $1$. The additional gauge sector helps in achieving a strong FOPT. We also consider $\chi,\Phi_{2}$ to be charged under $U(1)_{D}$ with equal charge $1$. This not only helps in stabilising DM, but also has interesting consequences for DM phenomenology to be discussed later. We ignore the couplings of singlet scalar with the standard model (SM) Higgs $H$ so that they do not affect the FOPT.
We calculate the complete potential including the tree level potential $V_{\rm tree}$, one-loop Coleman-Weinberg potential $V_{\rm CW}$oleman:1973jx along with the finite-temperature potential $V_{\rm th}$ Dolan:1973qd ; Quiros:1999jp . The thermal field-dependent masses of different particles coupled to the singlet scalar $\Phi_{1}$ are incorporated in the full potential. The zero-temperature masses of DM and $\Phi_{2}$ are denoted by $m_{\chi}$ and $m_{\Phi_{2}}$ respectively. We then calculate the critical temperature $T_{c}$ at which the potential in the $\Phi_{1}$ direction acquires another degenerate minima at $v_{c}=\phi_{1}(T=T_{c})$. The order parameter of the FOPT is defined as $v_{c}/T_{c}$ such that a stronger FOPT corresponds to a larger $v_{c}/T_{c}$. The FOPT then proceeds via tunneling, the rate of which is estimated by calculating the bounce action $S_{3}$ using the prescription in Linde:1980tt ; Adams:1993zs . The nucleation temperature $T_{n}$ is then calculated by comparing the tunneling rate with the Hubble expansion rate of the Universe $\Gamma(T_{n})=\mathcal{H}^{4}(T_{n})$. We also estimate the reheat temperature $T_{\rm RH}$ at the end of the FOPT due to the release of radiation energy. It is defined as $T_{\rm RH}={\rm Max}[T_{n},T_{\rm inf}]$ Baldes:2021vyz where $T_{\rm inf}$ is determined by equating density of radiation energy to that of energy released from the FOPT or equivalently $\Delta V_{\rm eff}$. The details of the FOPT is given in appendix A.
@@ III Co-genesis of Large neutrino asymmetry and Dark Matter
In order to calculate the final neutrino asymmetry and DM abundance starting with an initial dark asymmetry, we write the relevant Boltzmann equations for $\chi,\nu$ and their anti-particles $\bar{\chi},\bar{\nu}$ together with the one for $\Phi_{2}$. We consider temperature dependent masses of relevant particles such that $\chi,\bar{\chi}$ decay is allowed only for a finite window $T_{s}>T>T_{n}$. While neutrinos do not receive much thermal corrections below the electroweak scale, $\Phi_{2},\chi$ thermal masses play crucial role in determining the final asymmetries. The relevant Boltzmann equations for generating a large neutrino asymmetry from an initial dark sector asymmetry can be written as follows.
@@
$\frac{dY_{\chi}}{dz}$ $=-\frac{1}{\mathcal{H}z}\langle\Gamma_{\chi}\rangle\Theta(m_{\chi}-m_{\Phi_{2}})\left(Y_{\chi}-Y_{\chi}^{\rm eq}\frac{Y_{\Phi_{2}}}{Y_{\Phi_{2}}^{\rm eq}}\right)+\frac{1}{2\mathcal{H}z}\langle\Gamma_{\Phi_{2}}\rangle\Theta(m_{\Phi_{2}}-m_{\chi})\left(Y_{\Phi_{2}}-Y_{\Phi_{2}}^{\rm eq}\frac{Y_{\chi}}{Y_{\chi}^{\rm eq}}\right)$
@@
$-\frac{s}{\mathcal{H}z}\langle\sigma v_{\chi\bar{\chi}\rightarrow Z_{D}Z_{D}}\rangle\left(Y_{\chi}Y_{\bar{\chi}}-Y_{\chi}^{\rm eq}Y_{\bar{\chi}}^{\rm eq}\right),$
@@
$\frac{dY_{\bar{\chi}}}{dz}$ $=-\frac{1}{\mathcal{H}z}\langle\Gamma_{\bar{\chi}}\rangle\Theta(m_{\chi}-m_{\Phi_{2}})\left(Y_{\bar{\chi}}-Y_{\bar{\chi}}^{\rm eq}\frac{Y_{\Phi_{2}}}{Y_{\Phi_{2}}^{\rm eq}}\right)+\frac{1}{2\mathcal{H}z}\langle\Gamma_{\Phi_{2}}\rangle\Theta(m_{\Phi_{2}}-m_{\chi})\left(Y_{\Phi_{2}}-Y_{\Phi_{2}}^{\rm eq}\frac{Y_{\bar{\chi}}}{Y_{\bar{\chi}}^{\rm eq}}\right)$
@@
$-\frac{s}{\mathcal{H}z}\langle\sigma v_{\chi\bar{\chi}\rightarrow Z_{D}Z_{D}}\rangle\left(Y_{\chi}Y_{\bar{\chi}}-Y_{\chi}^{\rm eq}Y_{\bar{\chi}}^{\rm eq}\right),$
@@
$\frac{dY_{\Phi_{2}}}{dz}$ $=\frac{1}{\mathcal{H}z}\langle\Gamma_{\chi}\rangle\Theta(m_{\chi}-m_{\Phi_{2}})\left(Y_{\chi}-Y_{\chi}^{\rm eq}\frac{Y_{\Phi_{2}}}{Y_{\Phi_{2}}^{\rm eq}}\right)-\frac{1}{2\mathcal{H}z}\langle\Gamma_{\Phi_{2}}\rangle\Theta(m_{\Phi_{2}}-m_{\chi})\left(Y_{\Phi_{2}}-Y_{\Phi_{2}}^{\rm eq}\frac{Y_{\chi}}{Y_{\chi}^{\rm eq}}\right)$
@@
$+\frac{1}{\mathcal{H}z}\langle\Gamma_{\bar{\chi}}\rangle\Theta(m_{\chi}-m_{\Phi_{2}})\left(Y_{\bar{\chi}}-Y_{\bar{\chi}}^{\rm eq}\frac{Y_{\Phi_{2}}}{Y_{\Phi_{2}}^{\rm eq}}\right)-\frac{1}{2\mathcal{H}z}\langle\Gamma_{\Phi_{2}}\rangle\Theta(m_{\Phi_{2}}-m_{\chi})\left(Y_{\Phi_{2}}-Y_{\Phi_{2}}^{\rm eq}\frac{Y_{\bar{\chi}}}{Y_{\bar{\chi}}^{\rm eq}}\right),$
@@
$\frac{dY_{\nu}}{dz}=\frac{1}{\mathcal{H}z}\langle\Gamma_{\bar{\chi}}\rangle\Theta(m_{\chi}-m_{\Phi_{2}})\left(Y_{\bar{\chi}}-Y_{\bar{\chi}}^{\rm eq}\frac{Y_{\Phi_{2}}}{Y_{\Phi_{2}}^{\rm eq}}\right)+\frac{1}{2\mathcal{H}z}\langle\Gamma_{\Phi_{2}}\rangle\Theta(m_{\Phi_{2}}-m_{\chi})\left(Y_{\Phi_{2}}-Y_{\Phi_{2}}^{\rm eq}\frac{Y_{\bar{\chi}}}{Y_{\bar{\chi}}^{\rm eq}}\right),$
@@
$\frac{dY_{\bar{\nu}}}{dz}=\frac{1}{\mathcal{H}z}\langle\Gamma_{\chi}\rangle\Theta(m_{\chi}-m_{\Phi_{2}})\left(Y_{\chi}-Y_{\chi}^{\rm eq}\frac{Y_{\Phi_{2}}}{Y_{\Phi_{2}}^{\rm eq}}\right)+\frac{1}{2\mathcal{H}z}\langle\Gamma_{\Phi_{2}}\rangle\Theta(m_{\Phi_{2}}-m_{\chi})\left(Y_{\Phi_{2}}-Y_{\Phi_{2}}^{\rm eq}\frac{Y_{\chi}}{Y_{\chi}^{\rm eq}}\right).$ (4)
@@
Here $Y_{i}=n_{i}/s$ denotes comoving number density of species “i” with $s$ being the entropy density. Hubble expansion rate is denoted by $\mathcal{H}$ while the variable $z$ is $m_{\chi}/T$. The thermal averaged decay rate Kolb:1979qa ; Buchmuller:2004nz for $\chi\rightarrow\Phi_{2}\nu$ (and its conjugate) can be written as $\langle\Gamma_{\chi}\rangle=\frac{y^{2}_{1}}{32\pi}\left(1-\frac{m_{\Phi_{2}}^{2}}{m_{\chi}^{2}}\right)^{2}\frac{K_{1}(m_{\chi}/T)}{K_{2}(m_{\chi}/T)}$ where $K_{1,2}$ are modified Bessel functions of 1st, 2nd kind respectively. Similarly, $\langle\Gamma_{\Phi_{2}}\rangle=\frac{y^{2}_{1}}{16\pi}\left(1-\frac{m_{\chi_{2}}^{2}}{m_{\Phi_{2}}^{2}}\right)^{2}\frac{K_{1}(m_{\Phi_{2}}/T)}{K_{2}(m_{\Phi_{2}}/T)}$. The $\Theta$ functions associated with these decay widths ensures the kinematical constraints. The symmetric part of $\chi$ can annihilate into a pair of light $U(1)_{D}$ gauge bosons $Z_{D}$ with $\langle\sigma v_{\chi\bar{\chi}\rightarrow Z_{D}Z_{D}}\rangle$ denoting the corresponding thermal-averaged cross-section Gondolo:1990dk .
@@
The initial dark sector asymmetry $Y^{\rm in}_{\Delta\chi}$ in required amount can be generated in a variety of ways depending upon the UV completion discussed in VI. Also, this initial DM asymmetry can be related to the origin of baryon asymmetry in the spirit of cogenesis within the framework of asymmetric dark matter (ADM) Nussinov:1985xr ; Kaplan:2009ag paradigm [^3] See Petraki:2013wwa ; Zurek:2013wia for reviews of ADM scenarios..

### v2

# Large neutrino asymmetry from forbidden decay of dark matter

@@ Abstract
[…]
@@ I Introduction
If a mechanism to explain BAU also leaves a large neutrino asymmetry at low temperatures, it can have interesting implications for observations related to BBN and CMB. Although the observed BAU restricts the net lepton asymmetry around sphaleron decoupling temperature $(T_{\rm sph})$ to be of same order as $\eta_{B}$, it is possible to have a large lepton asymmetry at post-sphaleron epoch $(T<T_{\rm sph})$ if the asymmetry is stored in neutrinos [^1] Charge neutrality of the early Universe restricts the charge lepton asymmetry to be $\leq\eta_{B}$.. Such large neutrino asymmetry can enhance the effective relativistic degrees of freedom $N_{\rm eff}$ which can be measured at CMB experiments Escudero:2022okz . Additionally, a large neutrino asymmetry can also alter the predictions of BBN with observable consequences Mangano:2010ei ; Mangano:2011ip ; Castorina:2012md ; Escudero:2022okz ; Froustey:2024mgf . A large neutrino asymmetry also allows resonant production of sterile neutrino DM via Shi-Fuller mechanism Shi:1998km  [^2] One may refer to a review Drewes:2016upu for details of these mechanisms and relevant experimental constraints.. Alternately, such asymmetries can also alter the cosmological bounds on light sterile neutrinos Mirizzi:2012we . Large neutrino asymmetry can also have other consequences for gravitational wave by turning the QCD phase transition to first order Schwarz:2009ii , baryogenesis, Barrie:2017mmr ; Gao:2024fhm , astrophysical structure formation Zeng:2018pcv etc.
@@ II The framework
$\mathcal{L}\supset\big(y_{1}\overline{\chi}\Phi_{2}\nu+{\rm h.c.}\big)+m_{0}\bar{\chi}\chi+y_{\chi\Phi_{1}}\bar{\chi}\Phi_{1}\chi+V(\Phi_{1},\Phi_{2})$ (2)
@@
$V_{\rm tree}\equiv V(\Phi_{1},\Phi_{2})$ $=\frac{\lambda_{1}}{4}\left(\Phi_{1}^{2}-\frac{v_{D}^{2}}{2}\right)^{2}+\mu_{\Phi_{2}}^{2}|\Phi_{2}|^{2}+\lambda_{2}|\Phi_{2}|^{4}+\frac{\lambda_{\Phi_{1}\Phi_{2}}}{2}\Phi_{1}^{2}|\Phi_{2}|^{2}$ (3)
In the above scalar potential $v_{D}$ is the vacuum expectation value (VEV) acquired by the real scalar field $\Phi_{1}$ while driving the FOPT. It is worth mentioning that terms with odd powers of $\Phi_{1}$ can also be there in the above scalar potential as they are allowed by the chosen symmetries of the model. We have ignored them for simplicity as it keeps the number of free parameters minimal. The terms considered in the scalar potential of Eq. (3) are sufficient to induce a first-order phase transition and ensure forbidden decay of dark matter $\chi$ due to sudden change of $\Phi_{2}$ mass at nucleation temperature, the details of which are discussed below. A dark global $U(1)_{D}$ symmetry is considered under which $\Phi_{1}$ transforms trivially whereas $\chi,\Phi_{2}$ have $U(1)_{D}$ charges of $1$ each.
We calculate the complete potential including the tree level potential $V_{\rm tree}$, one-loop Coleman-Weinberg potential $V_{\rm CW}$ Coleman:1973jx along with the finite-temperature potential $V_{\rm th}$ Dolan:1973qd ; Quiros:1999jp . The thermal field-dependent masses of different particles coupled to the singlet scalar $\Phi_{1}$ are incorporated in the full potential. The zero-temperature masses of DM and $\Phi_{2}$ are denoted by $m_{\chi}$ and $m_{\Phi_{2}}$ respectively. We then calculate the critical temperature $T_{c}$ at which the potential in the $\Phi_{1}$ direction acquires another degenerate minima at $v_{c}=\phi_{1}(T=T_{c})$. The order parameter of the FOPT is defined as $v_{c}/T_{c}$ such that a stronger FOPT corresponds to a larger $v_{c}/T_{c}$. The FOPT then proceeds via tunneling, the rate of which is estimated by calculating the bounce action $S_{3}$ using the prescription in Linde:1980tt ; Adams:1993zs . The nucleation temperature $T_{n}$ is then calculated by comparing the tunneling rate with the Hubble expansion rate of the Universe $\Gamma(T_{n})=\mathcal{H}^{4}(T_{n})$. We also estimate the reheat temperature $T_{\rm RH}$ at the end of the FOPT due to the release of radiation energy. It is defined as $T_{\rm RH}={\rm Max}[T_{n},T_{\rm inf}]$ Baldes:2021vyz where $T_{\rm inf}$ is determined by equating density of radiation energy to that of energy released from the FOPT or equivalently $\Delta V_{\rm eff}$. The details of the FOPT is given in Appendix A.
@@ III Co-genesis of Large neutrino asymmetry and Dark Matter
In order to calculate the final neutrino asymmetry and DM abundance starting with an initial dark asymmetry, we write the relevant Boltzmann equations for $\chi,\Phi_{2},\nu$ and their anti-particles $\bar{\chi},\Phi_{2}^{\dagger},\bar{\nu}$. We consider temperature dependent masses of relevant particles such that $\chi,\bar{\chi}$ decay is allowed only for a finite window $T_{s}>T>T_{n}$. While neutrinos do not receive much thermal corrections below the electroweak scale, the thermal mass of $\Phi_{2}$ plays crucial role in determining the final asymmetries. The relevant Boltzmann equations for generating a large neutrino asymmetry from an initial dark sector asymmetry can be written as follows.
@@
$\frac{dY_{\chi}}{dz}$ $=-\frac{1}{\mathcal{H}z}\langle\Gamma_{\chi}\rangle\Theta(m_{\chi}-m_{\Phi_{2}})\left(Y_{\chi}-Y_{\chi}^{\rm eq}\frac{Y_{\Phi_{2}}}{Y_{\Phi_{2}}^{\rm eq}}\right)+\frac{1}{\mathcal{H}z}\langle\Gamma_{\Phi_{2}}\rangle\Theta(m_{\Phi_{2}}-m_{\chi})\left(Y_{\Phi_{2}}-Y_{\Phi_{2}}^{\rm eq}\frac{Y_{\chi}}{Y_{\chi}^{\rm eq}}\right)$
@@
$-\frac{s}{\mathcal{H}z}\langle\sigma v_{\chi\bar{\chi}\rightarrow\Phi_{1}\Phi_{1}}\rangle\left(Y_{\chi}Y_{\bar{\chi}}-Y_{\chi}^{\rm eq}Y_{\bar{\chi}}^{\rm eq}\right),$
@@
$\frac{dY_{\bar{\chi}}}{dz}$ $=-\frac{1}{\mathcal{H}z}\langle\Gamma_{\bar{\chi}}\rangle\Theta(m_{\chi}-m_{\Phi_{2}})\left(Y_{\bar{\chi}}-Y_{\bar{\chi}}^{\rm eq}\frac{Y_{\Phi_{2}^{\dagger}}}{Y_{\Phi_{2}^{\dagger}}^{\rm eq}}\right)+\frac{1}{\mathcal{H}z}\langle\Gamma_{\Phi_{2}^{\dagger}}\rangle\Theta(m_{\Phi_{2}}-m_{\chi})\left(Y_{\Phi_{2}^{\dagger}}-Y_{\Phi_{2}^{\dagger}}^{\rm eq}\frac{Y_{\bar{\chi}}}{Y_{\bar{\chi}}^{\rm eq}}\right)$
@@
$-\frac{s}{\mathcal{H}z}\langle\sigma v_{\chi\bar{\chi}\rightarrow\Phi_{1}\Phi_{1}}\rangle\left(Y_{\chi}Y_{\bar{\chi}}-Y_{\chi}^{\rm eq}Y_{\bar{\chi}}^{\rm eq}\right),$
@@
$\frac{dY_{\Phi_{2}}}{dz}$ $=\frac{1}{\mathcal{H}z}\langle\Gamma_{\chi}\rangle\Theta(m_{\chi}-m_{\Phi_{2}})\left(Y_{\chi}-Y_{\chi}^{\rm eq}\frac{Y_{\Phi_{2}}}{Y_{\Phi_{2}}^{\rm eq}}\right)-\frac{1}{\mathcal{H}z}\langle\Gamma_{\Phi_{2}}\rangle\Theta(m_{\Phi_{2}}-m_{\chi})\left(Y_{\Phi_{2}}-Y_{\Phi_{2}}^{\rm eq}\frac{Y_{\chi}}{Y_{\chi}^{\rm eq}}\right)$
@@
$-\frac{s}{\mathcal{H}z}\langle\sigma v_{\Phi_{2}\Phi_{2}^{\dagger}\rightarrow\Phi_{1}\Phi_{1}}\rangle\left(Y_{\Phi_{2}}Y_{\Phi_{2}^{\dagger}}-Y_{\Phi_{2}}^{\rm eq}Y_{\Phi_{2}^{\dagger}}^{\rm eq}\right)-\frac{s}{\mathcal{H}z}\langle\sigma v_{\Phi_{1}\Phi_{2}\rightarrow\Phi_{1}\Phi_{2}^{\dagger}}\rangle Y_{\Phi_{1}}^{\rm eq}\left(Y_{\Phi_{2}}-Y_{\Phi_{2}^{\dagger}}\right),$
@@
$\frac{dY_{\Phi_{2}^{\dagger}}}{dz}$ $=\frac{1}{\mathcal{H}z}\langle\Gamma_{\bar{\chi}}\rangle\Theta(m_{\chi}-m_{\Phi_{2}})\left(Y_{\bar{\chi}}-Y_{\bar{\chi}}^{\rm eq}\frac{Y_{\Phi_{2}^{\dagger}}}{Y_{\Phi_{2}^{\dagger}}^{\rm eq}}\right)-\frac{1}{\mathcal{H}z}\langle\Gamma_{\Phi_{2}^{\dagger}}\rangle\Theta(m_{\Phi_{2}}-m_{\chi})\left(Y_{\Phi_{2}^{\dagger}}-Y_{\Phi_{2}^{\dagger}}^{\rm eq}\frac{Y_{\bar{\chi}}}{Y_{\bar{\chi}}^{\rm eq}}\right)$
@@
$-\frac{s}{\mathcal{H}z}\langle\sigma v_{\Phi_{2}\Phi_{2}^{\dagger}\rightarrow\Phi_{1}\Phi_{1}}\rangle\left(Y_{\Phi_{2}}Y_{\Phi_{2}^{\dagger}}-Y_{\Phi_{2}}^{\rm eq}Y_{\Phi_{2}^{\dagger}}^{\rm eq}\right)+\frac{s}{\mathcal{H}z}\langle\sigma v_{\Phi_{1}\Phi_{2}\rightarrow\Phi_{1}\Phi_{2}^{\dagger}}\rangle Y_{\Phi_{1}}^{\rm eq}\left(Y_{\Phi_{2}}-Y_{\Phi_{2}^{\dagger}}\right),$
$\frac{dY_{\nu}}{dz}=\frac{1}{\mathcal{H}z}\langle\Gamma_{\chi}\rangle\Theta(m_{\chi}-m_{\Phi_{2}})\left(Y_{\chi}-Y_{\chi}^{\rm eq}\frac{Y_{\Phi_{2}}}{Y_{\Phi_{2}}^{\rm eq}}\right)+\frac{1}{\mathcal{H}z}\langle\Gamma_{\Phi_{2}^{\dagger}}\rangle\Theta(m_{\Phi_{2}}-m_{\chi})\left(Y_{\Phi_{2}^{\dagger}}-Y_{\Phi_{2}^{\dagger}}^{\rm eq}\frac{Y_{\bar{\chi}}}{Y_{\bar{\chi}}^{\rm eq}}\right),$
$\frac{dY_{\bar{\nu}}}{dz}=\frac{1}{\mathcal{H}z}\langle\Gamma_{\bar{\chi}}\rangle\Theta(m_{\chi}-m_{\Phi_{2}})\left(Y_{\bar{\chi}}-Y_{\bar{\chi}}^{\rm eq}\frac{Y_{\Phi_{2}^{\dagger}}}{Y_{\Phi_{2}^{\dagger}}^{\rm eq}}\right)+\frac{1}{\mathcal{H}z}\langle\Gamma_{\Phi_{2}}\rangle\Theta(m_{\Phi_{2}}-m_{\chi})\left(Y_{\Phi_{2}}-Y_{\Phi_{2}}^{\rm eq}\frac{Y_{\chi}}{Y_{\chi}^{\rm eq}}\right).$ (4)
@@
Here $Y_{i}=n_{i}/s$ denotes comoving number density of species “i” with $s$ being the entropy density. Hubble expansion rate is denoted by $\mathcal{H}$ while the variable $z$ is $m_{\chi}/T$. The thermal averaged decay rate Kolb:1979qa ; Buchmuller:2004nz for $\chi\rightarrow\Phi_{2}\nu$ (and its conjugate) can be written as $\langle\Gamma_{\chi}\rangle=\frac{y^{2}_{1}}{32\pi}m_{\chi}\left(1-\frac{m_{\Phi_{2}}^{2}}{m_{\chi}^{2}}\right)^{2}\frac{K_{1}(m_{\chi}/T)}{K_{2}(m_{\chi}/T)}$ where $K_{1,2}$ are modified Bessel functions of 1st, 2nd kind respectively. Similarly, $\langle\Gamma_{\Phi_{2}}\rangle=\frac{y^{2}_{1}}{16\pi}m_{\Phi_{2}}\left(1-\frac{m_{\chi}^{2}}{m_{\Phi_{2}}^{2}}\right)^{2}\frac{K_{1}(m_{\Phi_{2}}/T)}{K_{2}(m_{\Phi_{2}}/T)}$. The $\Theta$ functions associated with these decay widths ensures the kinematical constraints. The equations for $\nu$ and $\bar{\nu}$ should also have collision terms corresponding to the SM gauge interactions on the right hand side. These interactions are responsible for keeping the neutrino number density close to their equilibrium values. However, we have ignored those terms as they do not contribute to the neutrino asymmetries. The symmetric part of $\chi$ can annihilate into a pair of light singlet scalars $\Phi_{1}$ with $\langle\sigma v_{\chi\bar{\chi}\rightarrow\Phi_{1}\Phi_{1}}\rangle$ denoting the corresponding thermal-averaged cross-section Gondolo:1990dk . A symmetry-breaking term of the form $\lambda_{12}(\Phi_{2}\Phi_{2}\Phi_{1}\Phi_{1}+\text{h.c.})$ is also introduced, ensuring that the asymmetry accumulated in the complex scalar $\Phi_{2}$ gets washed out. While other explicit $U(1)_{D}$ breaking terms like $\mu_{12}\Phi_{2}\Phi^{2}_{1}+\mu^{\prime}_{12}\Phi^{2}_{2}\Phi_{1}$ can also cause such washout of $\Phi_{2}$ asymmetry, we consider only the quartic term in numerical calculations for illustrative purpose.
@@
The initial dark sector asymmetry $Y^{\rm in}_{\Delta\chi}$ in required amount can be generated in a variety of ways depending upon the UV completion discussed in section VI. Also, this initial DM asymmetry can be related to the origin of baryon asymmetry in the spirit of cogenesis within the framework of asymmetric dark matter (ADM) Nussinov:1985xr ; Kaplan:2009ag paradigm [^3] See Petraki:2013wwa ; Zurek:2013wia for reviews of ADM scenarios..